A paraplegic patient is stretching the hamstring m

Please post the steps to get the answer.

Thanks!
Answer #1

torque M on the hip = (r1*F1*sin(90))+(r2*F2*sin(90))


M = (0.2*400)+(0.5*80) = 120 N-m